In early 1999 an investor bought 1000 shares of skandia for K127 per share. During the year dividends were received at K1.15 per share. What is the investor's total return in K? Capital gains, dividend yield, and total return in percentage terms?
1
Expert's answer
2021-09-17T15:32:06-0400

Given Information:

Purchased 1000 shares foe K127 per share

Dividend received at K1.15 per shares

assuming, Selling price per share is K180



Part (a)

Amount spent to buy 1000 Shares "=1000 \times K127= K127,000"

Dividend Received "= 1000 \times K1.15=
 K1150"

Amount received from selling 1000 shares "= 1000 \times K180=

 K180,000"

Total return = Amount received from selling 1000 shares + Dividend Received - Amount spent to buy 1000 shares

"Total\space return = K180,000 + K1150 - K127,000

= K54,150"



Part (b)

Capital gains (in %) = [(Selling price of share - Buying price of share) "\div" Buying price of share]"\times" 100

= Capital gains (in %) "= [\frac{(K180 - K127) }{ K127}]\times 100= 41.73\%"


Dividend yield (in %) = (Dividend rate "\div" Buying price of share)"\times" 100

=Dividend yield (in%) "= (\frac{K1.15 }{K127})\times 100

 = 0.905\%"


Total return (in %) = [(Selling price of share + Dividend rate - Buying price of share) "\div" Buying price of share]"\times" 100

= Total return (in %) "= [\frac{(K180 + K1.15 - K127) }{ K127}]\times100= 42.63\%"